OpenStudy (anonymous):

the sum of the series 20c0 - 20c1 + 20c2 - 20c3+........+20c10 is

OpenStudy (amistre64):

20(c0-c1+c2-c3 + ... + c10); what are the coeefs?

OpenStudy (anonymous):

i think it should use binomial form (1-x)^20 or somewhat like it...

OpenStudy (anonymous):

no no no hold on.. its a combinatorics problem. the 'c' here means that

OpenStudy (anonymous):

use (1-x)^20+(1+x)^20 you will get the ans...

OpenStudy (anonymous):

deeprony have u understood the p problem??????

OpenStudy (amistre64):

20!/20! - 20!/19! + 20!/18!2! .... like that?

OpenStudy (anonymous):

"(1-x)^20+(1+x)^20" what is x here?? yea i understood that quetion=)

OpenStudy (anonymous):

yes mistre thats correct

OpenStudy (amistre64):

i stil gotta do these the brute force way ;)

OpenStudy (anonymous):

isint there a formula??

OpenStudy (anonymous):

(1-x)^20=20c0-20c1x+.......... now 20cn=20c20-n. and thus 20c20=20c0. and so on put x=1. 20c0-20c1+.....=0 2(20c0-20c1-.......+20c10)-20c10=0 hence ans is (20c10)/2

OpenStudy (anonymous):

can you explain what you did hre.. 20c0-20c1+.....=0 2(20c0-20c1-.......+20c10)-20c10=0

OpenStudy (anonymous):

i put just x=1 in the 1st expansion of (1-x)^20

OpenStudy (anonymous):

yea ok wat did u do thn??

OpenStudy (anonymous):

that is the way of doing.....using a binomial series....

OpenStudy (anonymous):

i had (1-x)^20=20c0-20c1x+20c2x^2.... put x=1 0=20c0-20c1+20c2......

OpenStudy (anonymous):

use 20c20=20c0,20c19=20c1,.......20c11=20c9.

OpenStudy (anonymous):

i got that olryt... i didnt understnd what u did xt

OpenStudy (anonymous):

*nxt

OpenStudy (anonymous):

you get the term 20c0,20c1,....20c9 twice but 20c10 only one...

OpenStudy (anonymous):

ahh i c... got it now...=))

OpenStudy (anonymous):

thus i write .. 0=2(20c0-20c1+20c2........+20c10)-20c10

OpenStudy (anonymous):

okkkkkkkkkkkkkkkkkk?????????

OpenStudy (anonymous):

yea!

OpenStudy (anonymous):

these are some tricks to do it...............

OpenStudy (anonymous):

is the answer 0